The question is......Two jets leave at the same time,one is flying eat, the other west. the speed of the jet that is flying west is 75km/h faster than the speed of the jet flying east. in 3 hr they are 3075km apart. let w represent the speed of the plane travelling west and e represent the speed of plane travelling east. you need to know the speed of each plane. write a system of equation to model this situation
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16949)

Plane going East x mph
Plane going West x+ 75 mph
..
They are moving away from each other

So add up their speed.

combined speed = x+x + 75
( 2 x + 75 )
Time = 3 hours
Distance = 3075 miles
Distance = speed * time

(2x+75)*3=3075
6.00 x + 225 = 3075
6 x = 3075 + -225
6 x = 2850
/ 6
x= 475 mph
Plane going East = 475 mph
Plane going West 475 + 75 = 550 mph
m.ananth@hotmail.ca

Answer by MathTherapy(10863)

The question is......Two jets leave at the same time,one is flying eat, the other west. the speed of the jet that is flying
west is 75km/h faster than the speed of the jet flying east. in 3 hr they are 3075km apart. let w represent the speed of the
plane travelling west and e represent the speed of plane travelling east. you need to know the speed of each plane. write a
system of equation to model this situation
******************************************
The other "PERSON" totally ignored the given instructions, regarding the assigned variables. It/He/She even came up with a solution,
although one wasn't requested. 

As requested, speeds of westbound and eastbound jets are w, and e, respectively
Since the westbound jet is travelling 75 km/h faster than the eastbound, we get the following SPEED equation: w = e + 75 ---- eq (i)

Also, it takes the jets 3 hrs to be 3,075 km apart, at which time, the westbound had travelled "3w" km, and the eastbound, "3e" km,
resulting in the following DISTANCE equation: 3w + 3e = 3,075
                                                                               3(w + e) = 3(1,025)
                                                                                    w + e = 1,025 --- eq (ii)
                                                                                           w = e + 75 -- eq (i)

The above two-equation system can be easily solved using SUBSTITUTION.
It can also be solved using ELIMINATION, but written as follows: w - e = 75 ----- eq (i)
                                                                                                               w + e = 1,025 -- eq (ii)
         
That's IT!. That's ALL that's needed.